find the equation of the line through (2,3) and (-1,2)
Answer by unit(6)

find the equation of the line through (2,3) and (-1,2)
Use equation of a line through two points
y -y1 = m(x-x1), where m = (y2-y1)/(x2-x1)
m= (2-3)/(-1-2) = -1/-3 = 1/3
Equation is
y-3 = 1/3(x-2)
3y-9=x-2
3y = x + 7